Determine the mass in grams of 15.0 moles of BaCl2
Answer #1

Molar mass of BaCl2,

MM = 1*MM(Ba) + 2*MM(Cl)

= 1*137.3 + 2*35.45

= 208.2 g/mol

use:

mass of BaCl2,

m = number of mol * molar mass

= 15 mol * 2.082*10^2 g/mol

= 3.123*10^3 g

Answer: 3.12*10^3 g
Answer #2

3.12*10^3 g